# Diffusion Models: A Comprehensive Survey of Methods and Applications

### v13

@@ 2.3. Stochastic Differential Equations (Score SDEs)
| Yang et al. (Yang et al., 2024f) | 2024 | Discrete | |||
@@ 4.1. Noise Schedule Optimization
Another line of works (Huang et al., 2024c; Yang et al., 2024f) propose to the modify diffusion trajectory through the integration of cross-modality information. Specifically, the cross-modal information, denoted as , is extracted from any conditional input and original sample with relational network . And then it can be injected to the forward process as an additional bias to adapt diffusion trajectory:
where is a non-negative scalar that control the magnitude of the bias term. It is important to note that with this modification, the forward process ceases to be a Markovian chain. ContextDiff (Yang et al., 2024f) introduces a general framework to jointly learn the cross-modal relational network and the diffusion model, and derives the VLB and sampling procedure for this modified diffusion process.
@@ 6. Connections with Other Generative Models
| Large Language Model | Zhang et al. (Zhang et al., 2024) | 2024 |
| Yang et al. (Yang et al., 2024e) | 2024 | |
| Yang et al. (Yang et al., 2024c) | 2024 | |
@@ 6.1. Large Language Models and Connections with Diffusion Models
Large Language Models (LLMs) (Brown et al., 2020; Anil et al., 2023; Achiam et al., 2023; Jiang et al., 2024; Yang et al., 2024d) have profoundly impacted the AI community, and showcased the advanced language comprehension and reasoning abilities. Recent works begin to extend their impressive reasoning abilities to visual generative tasks for overall generation planning. The collaboration between LLMs (ChatGPT, 2022; OpenAI, 2023b; Yang et al., 2024d) and diffusion models (Ramesh et al., 2022; Betker et al., 2023; Yang et al., 2023a, 2024f) can significantly improve the text-image alignment as well as the quality of generated images (Zhang et al., 2024; Lian et al., 2023; Tian et al., 2024). For instance, RealCompo (Zhang et al., 2024) utilizes LLMs to enhance the compositional generation of diffusion models by generating images grounded on bounding box layouts from the LLM. EditWorld (Yang et al., 2024e) composes a set of LLMs and pretrained diffusion models to generate an image editing dataset that contains numerous instructions with world knowledge (Ha and Schmidhuber, 2018). VideoTetris (Tian et al., 2024) uses the LLM to decompose text prompts along temporal axis for guiding video generation with smoother and more reasonable transitions. Further, RPG (Yang et al., 2024c) leverages the vision-language prior of multimodal LLMs to reason out complementary spatial layouts from text prompt, and manipulates the object compositions for diffusion models in both text-guided image generation and editing process, achieving SOTA performance in compositional synthesis scenarios.
@@ 6.3. Generative Adversarial Networks and Connections with Diffusion Models
One of the issues of GAN is the instability in the training process, which is mainly caused by the non-overlapping between the distribution of input data and that of the generated data. One solution is to inject noise into the discriminator input for widening the support of both the generator and discriminator distributions. Taking advantage of the flexible diffusion model, Wang et al. (2022) (Wang et al., 2022b) inject noise to the discriminator with an adaptive noise schedule determined by a diffusion model. On the other hand, GAN can facilitate sampling speed of diffusion models. Xiao et al. (2021) (Xiao et al., 2021) show that slow sampling is caused by the Gaussian assumption in the denoising step, which is justified only for small step sizes. As such, each denoising step is modeled by a conditional GAN, allowing larger step size.
@@ 6.4. Normalizing Flows and Connections with Diffusion Models
Similar to the continuous setting, normalizing flows allow for the retrieval of the exact log-likelihood through a change of variable formula. However, the bijection requirement limits the modeling of complex data in both practical and theoretical contexts (Wu et al., 2020a; Cornish et al., 2020). Several works attempt to relax this bijection requirement (Wu et al., 2020a; Dinh et al., 2019). For example, DiffFlow (Zhang and Chen, 2021) introduces a generative modeling algorithm that combines the benefits of both flow-based and diffusion models. As a result, DiffFlow produces sharper boundaries than normalizing flow and learns more general distributions with fewer discretization steps compared to diffusion probabilistic models.
Implicit Nonlinear Diffusion Model (INDM) (Kim et al., 2022c) optimizes the pre-encoding process of latent diffusion, which first encodes the original data into the latent space using normalizing flow, and then performs diffusion in the latent space. Using theory from stochastic calculus, transforming data using normalizing flow can be seen as learning a non-linear SDE which drift and diffusion coefficient are determined by the normalizing flow. The score matching objective in INDM equals to a combination of objectives in diffusion model and normalizing flow. Using a non-linear diffusion process, INDM can effectively improve the likelihood and the sampling speed.
@@ 6.6. Energy-based Models and Connections with Diffusion Models
Although EBMs have a number of desirable properties, two challenges remain for modeling high-dimensional data. First, learning EBMs by maximizing the likelihood requires MCMC method to generate samples from the model, which can be very computationally expensive. Second, as demonstrated in (Nijkamp et al., 2019a), the energy potentials learned with non-convergent MCMC are not stable, in the sense that samples from long-run Markov chains can be significantly different from the observed samples, and thus it is difficult to evaluate the learned energy potentials. In a recent study, Gao et al. (2021) (Gao et al., 2020b) present a diffusion recovery likelihood method to tractably learn samples from a sequence of EBMs in the reverse process of the diffusion model. Each EBM is trained with recovery likelihood, which aims to maximize the conditional probability of the data at a certain noise level, given their noisy versions at a higher noise level. EBMs maximize the recovery likelihood because it is more tractable than marginal likelihood, as sampling from the conditional distributions is much easier than sampling from the marginal distributions. This model can generate high-quality samples, and long-run MCMC samples from the conditional distributions still resemble realistic images.
@@ 7. Applications of Diffusion Models
| Super Resolution, Inpainting, Restoration, Translation, and Editing | (Song et al., 2021b),(Chung et al., 2022a),(Meng et al., 2021a),(Kawar et al., 2022a),(Yang et al., 2023a),(Yang et al., 2024c) | |
@@
| Video Generation | (Harvey et al., 2022),(Ho et al., 2022c),(Yang et al., 2022b),(Zhang et al., 2022a),(Singer et al., 2022),(Ho et al., 2022a),(Wu et al., 2022a),(Qi et al., 2023) | |
| Point Cloud Completion and Generation | (Zhou et al., 2021),(Luo and Hu, 2021a),(Lyu et al., 2021),(Liu et al., 2022),(Zeng et al., 2022) | |
| Generating Data from Diffusion Models | (Yang et al., 2024e),(Brooks et al., 2023),(Zhu et al., 2024) | |
@@
| Temporal Data Modeling | Time Series Imputation | (Tashiro et al., 2021),(Alcaraz and Strodthoff, 2022),(Park et al., 2021) |
| Time Series Forecasting | (Rasul et al., 2021b),(Alcaraz and Strodthoff, 2022) | |
@@
| Multi-Modal Learning | Text-to-Image Generation | (Avrahami et al., 2022),(Ramesh et al., 2022),(Saharia et al., 2022b),(Nichol et al., 2022),(Gu et al., 2022),(Ruiz et al., 2022),(Kawar et al., 2022c),(Xu et al., 2022b),(Zhang and Agrawala, 2023),(Yang et al., 2023a),(Yang et al., 2024f),(Yang et al., 2024c) |
@@
| Text-to-3D Generation | (Xu et al., 2022a),(Lin et al., 2022),(Poole et al., 2022),(Zeng et al., 2023) | |
@@
| Text-to-Video Generation | (Singer et al., 2022),(Ho et al., 2022a),(Wu et al., 2022a),(Qi et al., 2023),(Henschel et al., 2024),(Yang et al., 2024f),(Tian et al., 2024) | |
@@ 7.1.3. Condition Diffusion on Texts, Images, and Semantic Maps
Recent researches begin to condition diffusion process on the guidance of more semantics, such as texts, images, and semantic maps, to better express rich semantics in samples. DiffuSeq (Gong et al., 2023) conditions on texts and proposes a seq-to-seq diffusion framework that helps with four NLP tasks. SDEdit (Meng et al., 2021a) conditions on a styled images to make image-to-image translation, while LDM (Rombach et al., 2022) unifies these semantic conditions with flexible latent diffusion. Kindly note that if conditions and diffusion targets are of different modalities, pre-alignment (Yang et al., 2022a; Ramesh et al., 2022) is a practical way to strengthen the guided diffusion. unCLIP (Ramesh et al., 2022) and ConPreDiff (Yang et al., 2023a) leverage CLIP latents in text-to-image generation, which have align the semantics between images and texts. RPG (Yang et al., 2024c) conditions on complementary rectangle and contour regions to enable compositional text-to-image generation and complex text-guided image editing. ContextDiff (Yang et al., 2024f) proposes a universal forward-backward consistent diffusion model for better conditioning on various input modalities.
@@ 7.2.4. Generating Data from Diffusion Models
Synthesizing datasets from generative models can effectively advance various tasks like classification (Wang et al., 2022c; Bansal and Grover, 2023; Zhu et al., 2024). Recent works have begun to utilize diffusion models to achieve this goal for vision tasks. For example, Trabucco et al. (2023) adopt diffusion models to make effective data augmentation for few-shot image classification. DistDiff (Zhu et al., 2024) proposes a training-free data expansion framework with a distribution-aware diffusion model. It constructs hierarchical prototypes to approximate the real data distribution, and optimizes latent data points in generation process with hierarchical energy guidance. InstructPix2Pix (Brooks et al., 2023) leverages two large pretrained models (i.e., GPT-3 and Stable Diffusion) to generate a large dataset of input-goal-instruction triplet examples, and trains an instruction-following image editing model on the dataset. To enable image editing to reflect chalenging world knowledge and dynamics from both real physical world and virtual media, EditWorld (Yang et al., 2024e), introduces a new task named world-instructed image editing, as the data examples presented in Fig. 6. EditWorld proposes an innovative compositional framework with a set of pretrained LLMs and Diffusion Models, illustrated in Fig. 7, to synthesize a world-instructed training dataset for instruction-following image editing.
@@ 7.4.1. Text-to-Image Generation
Inspired by the ability of guided diffusion models (Dhariwal and Nichol, 2021; Ho and Salimans, 2022) to generate photorealistic samples and the ability of text-to-image models to handle free-form prompts, GLIDE (Nichol et al., 2022) applies guided diffusion to the application of text-conditioned image synthesis. VQ-Diffusion (Gu et al., 2022) proposes a vector-quantized diffusion model for text-to-image generation, and it eliminates the unidirectional bias and avoids accumulative prediction errors. Versatile Diffusion (Xu et al., 2022b) proposes the first unified multi-flow multimodal diffusion framework, which supports image-to-text, image-variation, text-to-image, and text-variation, and can be further extended to other applications such as semantic-style disentanglement, image-text dual-guided generation, latent image-to-text-to-image editing, and more. Following Versatile Diffusion, UniDiffuser (Bao et al., 2023) proposes a unified diffusion model framework based on Transformer, which can fit multimodal data distributions and simultaneously handle text-to-image, image-to-text, and joint image-text generation tasks. ConPreDiff (Yang et al., 2023a) for the first time incorporates context prediction into text-to-image diffusion models, and significantly improves generation performance without additional inference costs. ContextDiff (Yang et al., 2024f) proposes general contextualized diffusion model by incorporating the cross-modal context encompassing interactions and alignments into forward and reverse processes. A qualitative comparison between these models are presented in Fig. 9.
@@
A new interesting line of diffusion model research is to leverage the pre-trained text-to-image diffusion model for more complex or fine-grained control of synthesis results. DreamBooth (Ruiz et al., 2022) presents the first technique that tackles the new challenging problem of subject-driven generation, allowing users, from just a few casually captured images of a subject, to recontextualize subjects, modify their properties, original art renditions, and more. Different from those image diffusion models conditioned on text prompts, ControlNet (Zhang and Agrawala, 2023) attempts to control pre-trained large diffusion models to support additional semantic maps, like edge maps, segmentation maps, keypoints, shape normals, depths, etc. However, most methods often face challenges when handling complex text prompts involving multiple objects with multiple attributes and relationships. To this end, RPG (Yang et al., 2024c) proposes a brand new training-free text-to-image generation/editing framework harnessing the powerful chain-of-thought reasoning ability of multimodal LLMs (Zhang et al., 2023) to enhance the compositionality of text-to-image diffusion models. This new RPG framework unifies both text-guided image generation (in Fig. 10) and image editing (in Fig. 11) tasks in a closed-loop fashion. Notably, as demonstrated in Fig. 12, RPG outperforms all SOTA methods, such as SDXL (Podell et al., 2023) and DALL-E 3 (Betker et al., 2023), demonstrate its superiority. Furthermore, RPG framework is user-friendly, and can generalize to different MLLM architectures and diffusion backbones (e.g., ControlNet).

### v14

@@ 2.3. Stochastic Differential Equations (Score SDEs)
| Yang et al. (Yang et al., 2024h) | 2024 | Discrete | |||
@@ 4.1. Noise Schedule Optimization
Another line of works (Huang et al., 2024c; Yang et al., 2024h) propose to the modify diffusion trajectory through the integration of cross-modality information. Specifically, the cross-modal information, denoted as , is extracted from any conditional input and original sample with relational network . And then it can be injected to the forward process as an additional bias to adapt diffusion trajectory:
where is a non-negative scalar that control the magnitude of the bias term. It is important to note that with this modification, the forward process ceases to be a Markovian chain. ContextDiff (Yang et al., 2024h) introduces a general framework to jointly learn the cross-modal relational network and the diffusion model, and derives the VLB and sampling procedure for this modified diffusion process.
@@ 6. Connections with Other Generative Models
| Large Language Model | Zhang et al. (Zhang et al., 2024a) | 2024 |
| Yang et al. (Yang et al., 2024f) | 2024 | |
| Yang et al. (Yang et al., 2024d) | 2024 | |
@@ 6.1. Large Language Models and Connections with Diffusion Models
Large Language Models (LLMs) (Brown et al., 2020; Anil et al., 2023; Achiam et al., 2023; Jiang et al., 2024; Yang et al., 2024e) have profoundly impacted the AI community, and showcased the advanced language comprehension and reasoning abilities. Recent works begin to extend their impressive reasoning abilities to visual generative tasks for overall generation planning. The collaboration between LLMs (ChatGPT, 2022; OpenAI, 2023b; Yang et al., 2024e) and diffusion models (Ramesh et al., 2022; Betker et al., 2023; Yang et al., 2023a, 2024h) can significantly improve the text-image alignment as well as the quality of generated images (Zhang et al., 2024a; Lian et al., 2023; Tian et al., 2024). For instance, RealCompo (Zhang et al., 2024a) utilizes LLMs to enhance the compositional generation of diffusion models by generating images grounded on bounding box layouts from the LLM. EditWorld (Yang et al., 2024f) composes a set of LLMs and pretrained diffusion models to generate an image editing dataset that contains numerous instructions with world knowledge (Ha and Schmidhuber, 2018). VideoTetris (Tian et al., 2024) uses the LLM to decompose text prompts along temporal axis for guiding video generation with smoother and more reasonable transitions. SemanticSDS (Yang et al., 2024g) and Trans4D (Zeng et al., 2024) extend the planning ability of LLMs to facilitate more complex 3D and 4D diffusion generation. Notably, RPG (Yang et al., 2024d) leverages the vision-language prior of multimodal LLMs to reason out complementary spatial layouts from text prompt, and manipulates the object compositions for diffusion models in both text-guided image generation and editing process, achieving SOTA performance in compositional synthesis scenarios and providing guidance for subsequent research.
@@ 6.3. Generative Adversarial Networks and Connections with Diffusion Models
One of the issues of GAN is the instability in the training process, which is mainly caused by the non-overlapping between the distribution of input data and that of the generated data. One solution is to inject noise into the discriminator input for widening the support of both the generator and discriminator distributions. Taking advantage of the flexible diffusion model, Wang et al. (2022) (Wang et al., 2022b) inject noise to the discriminator with an adaptive noise schedule determined by a diffusion model. On the other hand, GAN can facilitate sampling speed of diffusion models. Xiao et al. (2021) (Xiao et al., 2021) show that slow sampling is caused by the Gaussian assumption in the denoising step, which is justified only for small step sizes. As such, each denoising step is modeled by a conditional GAN, allowing larger step size. To ensure the diffusion model captures authentic manifold structures in the data distribution, SADM (Yang et al., 2024b) advocates adversarial training of the diffusion generator against a novel structure discriminator in a minimax game, distinguishing real manifold structures from the generated ones.
@@ 6.4. Normalizing Flows and Connections with Diffusion Models
Similar to the continuous setting, normalizing flows allow for the retrieval of the exact log-likelihood through a change of variable formula. However, the bijection requirement limits the modeling of complex data in both practical and theoretical contexts (Wu et al., 2020a; Cornish et al., 2020). Several works attempt to relax this bijection requirement (Wu et al., 2020a; Dinh et al., 2019). For example, DiffFlow (Zhang and Chen, 2021) introduces a generative modeling algorithm that combines the benefits of both flow-based and diffusion models. As a result, DiffFlow produces sharper boundaries than normalizing flow and learns more general distributions with fewer discretization steps compared to diffusion probabilistic models. Implicit Nonlinear Diffusion Model (INDM) (Kim et al., 2022c) optimizes the pre-encoding process of latent diffusion, which first encodes the original data into the latent space using normalizing flow, and then performs diffusion in the latent space. Using a non-linear diffusion process, INDM can effectively improve the likelihood and the sampling speed.
To scale up the training of CNFs, recent works propose efficient simulation-free approaches (Lipman et al., 2022; Albergo and Vanden-Eijnden, 2022; Liu et al., 2022a) by parameterizing a vector field which flows from noise samples to data samples. Lipman et al. (2022) (Lipman et al., 2022) propose Flow Matching (FM) to train CNFs based on constructing explicit conditional probability paths between the noise distribution and each data sample. Wang et al. (2024) (Wang et al., 2024) conduct an in-depth analysis of the essence of rectification in rectified flow (Liu et al., 2022a) and extend it to rectified diffusion. Besides, they identify that it is not straightness but first-order property is the essential training target of rectified diffusion with theoretical derivations. Yang et al. (2024) further propose Consistency Flow Matching (Yang et al., 2024i), a novel FM method that explicitly enforces self-consistency in the velocity field. Consistency Flow Matching (Yang et al., 2024i) directly defines straight flows starting from different times to the same endpoint, imposing constraints on their velocity values:
where is the uniform distribution on , is a positive scalar, denotes a time interval which is a small and positive scalar. denotes the running average of past values of using exponential moving average (EMA), and follows a pre-defined distribution which can be efficiently sampled, for example, VP-SDE (Ho et al., 2020) or OT path (Lipman et al., 2022). In this way, Consistency Flow Matching (Yang et al., 2024i) innovatively bridges consistency models and flow matching models through the novel concept of straight flows characterized by velocity consistency.
@@ 6.6. Energy-based Models and Connections with Diffusion Models
Although EBMs have a number of desirable properties, two challenges remain for modeling high-dimensional data. First, learning EBMs by maximizing the likelihood requires MCMC method to generate samples from the model, which can be very computationally expensive. Second, as demonstrated in (Nijkamp et al., 2019a), the energy potentials learned with non-convergent MCMC are not stable, in the sense that samples from long-run Markov chains can be significantly different from the observed samples, and thus it is difficult to evaluate the learned energy potentials. In a recent study, Gao et al. (2021) (Gao et al., 2020b) present a diffusion recovery likelihood method to tractably learn samples from a sequence of EBMs in the reverse process of the diffusion model. Each EBM is trained with recovery likelihood, which aims to maximize the conditional probability of the data at a certain noise level, given their noisy versions at a higher noise level. EBMs maximize the recovery likelihood because it is more tractable than marginal likelihood, as sampling from the conditional distributions is much easier than sampling from the marginal distributions.
@@ 7. Applications of Diffusion Models
| Super Resolution, Inpainting, Restoration, Translation, and Editing | (Song et al., 2021b),(Chung et al., 2022a),(Meng et al., 2021a),(Kawar et al., 2022a),(Yang et al., 2023a),(Yang et al., 2024d) | |
@@
| Video Generation | (Harvey et al., 2022),(Ho et al., 2022c),(Yang et al., 2022b),(Zhang et al., 2022a),(Singer et al., 2022),(Ho et al., 2022a),(Wu et al., 2022a),(Qi et al., 2023),(Tian et al., 2024) | |
| Point Cloud Completion and Generation | (Zhou et al., 2021),(Luo and Hu, 2021a),(Lyu et al., 2021),(Liu et al., 2022b),(Zeng et al., 2022) | |
| Generating Data from Diffusion Models | (Yang et al., 2024f),(Brooks et al., 2023),(Zhu et al., 2024) | |
@@
| Temporal Data Modeling | Time Series Imputation | (Tashiro et al., 2021),(Alcaraz and Strodthoff, 2022),(Park et al., 2021),(Liu et al., 2024) |
| Time Series Forecasting | (Rasul et al., 2021b),(Alcaraz and Strodthoff, 2022),(Liu et al., 2024) | |
@@
| Multi-Modal Learning | Text-to-Image Generation | (Avrahami et al., 2022),(Ramesh et al., 2022),(Saharia et al., 2022b),(Nichol et al., 2022),(Gu et al., 2022),(Ruiz et al., 2022),(Kawar et al., 2022c),(Xu et al., 2022b),(Zhang and Agrawala, 2023),(Yang et al., 2023a),(Yang et al., 2024h),(Yang et al., 2024d),(Zhang et al., 2024b) |
@@
| Text-to-3D/4D Generation | (Xu et al., 2022a),(Lin et al., 2022),(Poole et al., 2022),(Zeng et al., 2023),(Yang et al., 2024g),(Zeng et al., 2024) | |
@@
| Text-to-Video Generation | (Singer et al., 2022),(Ho et al., 2022a),(Wu et al., 2022a),(Qi et al., 2023),(Henschel et al., 2024),(Yang et al., 2024h),(Tian et al., 2024) | |
@@ 7.1.1. Conditioning Mechanisms in Diffusion Models
7.1.2. Diffusion with DPO/RLHF
Building on the success of reinforcement learning from human feedback (RLHF) in Large Language Models (LLMs) (Ouyang et al., 2022; Bai et al., 2022), numerous methods in diffusion models have attempted to use similar approaches for model alignment (Lee et al., 2023; Fan et al., 2024). Some methods use a pretrained reward model or train a new one to guide the generation process. For instance, ImageReward (Xu et al., 2024) manually annotated a large dataset of human-preferred images and trained a reward model to assess the alignment between images and human preferences. Some methods bypass the training of a reward model and directly finetune diffusion models on human preference datasets (Yang et al., 2024c). Diffusion-DPO (Wallace et al., 2024) reformulates Direct Preference Optimization (DPO) to account for a diffusion model’s notion of likelihood, utilizing the evidence lower bound to derive a differentiable objective. Recently, Zhang et al. (2024) propose IterComp (Zhang et al., 2024b) to iteratively align the base diffusion model with composition-aware model preferences from the model gallery, consisting of six powerful open-source diffusion models, effectively enhancing the performance of base model on conditional diffusion generation. As demonstrated in Fig. 4, IterComp (Zhang et al., 2024b) outperforms other three types of conditional diffusion methods while achieving the best inference efficiency.
@@ 7.1.4. Condition Diffusion on Texts, Images, and Semantic Maps
Recent researches begin to condition diffusion process on the guidance of more semantics, such as texts, images, and semantic maps, to better express rich semantics in samples. DiffuSeq (Gong et al., 2023) conditions on texts and proposes a seq-to-seq diffusion framework that helps with four NLP tasks. SDEdit (Meng et al., 2021a) conditions on a styled images to make image-to-image translation, while LDM (Rombach et al., 2022) unifies these semantic conditions with flexible latent diffusion. Kindly note that if conditions and diffusion targets are of different modalities, pre-alignment (Yang et al., 2022a; Ramesh et al., 2022) is a practical way to strengthen the guided diffusion. unCLIP (Ramesh et al., 2022) and ConPreDiff (Yang et al., 2023a) leverage CLIP latents in text-to-image generation, which have align the semantics between images and texts. RPG (Yang et al., 2024d) conditions on complementary rectangle and contour regions to enable compositional text-to-image generation and complex text-guided image editing. ContextDiff (Yang et al., 2024h) proposes a universal forward-backward consistent diffusion model for better conditioning on various input modalities.
@@ 7.2.4. Generating Data from Diffusion Models
Synthesizing datasets from generative models can effectively advance various tasks like classification (Wang et al., 2022c; Bansal and Grover, 2023; Zhu et al., 2024). Recent works have begun to utilize diffusion models to achieve this goal for vision tasks. For example, Trabucco et al. (2023) adopt diffusion models to make effective data augmentation for few-shot image classification. DistDiff (Zhu et al., 2024) proposes a training-free data expansion framework with a distribution-aware diffusion model. It constructs hierarchical prototypes to approximate the real data distribution, and optimizes latent data points in generation process with hierarchical energy guidance. InstructPix2Pix (Brooks et al., 2023) leverages two large pretrained models (i.e., GPT-3 and Stable Diffusion) to generate a large dataset of input-goal-instruction triplet examples, and trains an instruction-following image editing model on the dataset. To enable image editing to reflect chalenging world knowledge and dynamics from both real physical world and virtual media, EditWorld (Yang et al., 2024f), introduces a new task named world-instructed image editing, as the data examples presented in Fig. 7. EditWorld proposes an innovative compositional framework with a set of pretrained LLMs and Diffusion Models, illustrated in Fig. 8, to synthesize a world-instructed training dataset for instruction-following image editing.
@@ 7.4.1. Text-to-Image Generation
Inspired by the ability of guided diffusion models (Dhariwal and Nichol, 2021; Ho and Salimans, 2022) to generate photorealistic samples and the ability of text-to-image models to handle free-form prompts, GLIDE (Nichol et al., 2022) applies guided diffusion to the application of text-conditioned image synthesis. VQ-Diffusion (Gu et al., 2022) proposes a vector-quantized diffusion model for text-to-image generation, and it eliminates the unidirectional bias and avoids accumulative prediction errors. Versatile Diffusion (Xu et al., 2022b) proposes the first unified multi-flow multimodal diffusion framework, which supports image-to-text, image-variation, text-to-image, and text-variation, and can be further extended to other applications such as semantic-style disentanglement, image-text dual-guided generation, latent image-to-text-to-image editing, and more. Following Versatile Diffusion, UniDiffuser (Bao et al., 2023) proposes a unified diffusion model framework based on Transformer, which can fit multimodal data distributions and simultaneously handle text-to-image, image-to-text, and joint image-text generation tasks. ConPreDiff (Yang et al., 2023a) for the first time incorporates context prediction into text-to-image diffusion models, and significantly improves generation performance without additional inference costs. ContextDiff (Yang et al., 2024h) proposes general contextualized diffusion model by incorporating the cross-modal context encompassing interactions and alignments into forward and reverse processes. A qualitative comparison between these models are presented in Fig. 10.
@@
A new interesting line of diffusion model research is to leverage the pre-trained text-to-image diffusion model for more complex or fine-grained control of synthesis results. DreamBooth (Ruiz et al., 2022) presents the first technique that tackles the new challenging problem of subject-driven generation, allowing users, from just a few casually captured images of a subject, to recontextualize subjects, modify their properties, original art renditions, and more. Different from those image diffusion models conditioned on text prompts, ControlNet (Zhang and Agrawala, 2023) attempts to control pre-trained large diffusion models to support additional semantic maps, like edge maps, segmentation maps, keypoints, shape normals, depths, etc. However, most methods often face challenges when handling complex text prompts involving multiple objects with multiple attributes and relationships. To this end, RPG (Yang et al., 2024d) proposes a brand new training-free text-to-image generation/editing framework harnessing the powerful chain-of-thought reasoning ability of multimodal LLMs (Zhang et al., 2023) to enhance the compositionality of text-to-image diffusion models. This new RPG framework unifies both text-guided image generation (in Fig. 11) and image editing (in Fig. 12) tasks in a closed-loop fashion. Notably, as demonstrated in Fig. 13, RPG outperforms all SOTA methods, such as SDXL (Podell et al., 2023) and DALL-E 3 (Betker et al., 2023), demonstrate its superiority. Furthermore, RPG framework is user-friendly, and can generalize to different MLLM architectures and diffusion backbones (e.g., ControlNet).